Electrolysis of Dilute sodium nitrate.
Q:Write the half-equation for anode and cathode reactions.
1
Expert's answer
2017-02-15T07:22:19-0500
Solution:
When diluted sodium nitrate will be electrolysis of water:
Then half-reaction at the cathode and anode will look like this:
At the cathode:
C (-): 2H20 + 2e = H2 + 2OH-
At the anode:
A (+): 2H2O - 4e = O2 + 4H+
Answer: C (-): 2H20 + 2e = H2 + 2OH- ; A (+): 2H2O - 4e = O2 + 4H+.